OpenStudy (anonymous):

How do you solve this? the square root of x squared + y squared is not equal to x + y and I need to provide a real value for each variable

OpenStudy (anonymous):

sqrt(x^2+y^2) not= x+y

Directrix (directrix):

>>need to provide a real value for each variable √ (3² + 4²) = 3 + 4 ? Yes or No? @needmathhelpplease